Solve the system with the elimination method.
4x-y=11
3x=3-y

Answer by rapaljer(4551)

4x-y=11
3x=3-y

In the second equation get it into standard form, by adding +y to each side:

4x-y=11
3x +y =3

Now, just add the two equations together:
7x = 14
x= 2

Next, substitute x =2 into EITHER equation. The first will do.
4x-y = 11
4(2)- y = 11
8-y = 11
-y = 3
y = -3

The point of solution is (2,-3).

Check by substituting both values into the OTHER equation:
3x= 3-y
3(2) = 3-(-3)
6=6
It checks!

R^2 at SCC